OpenStudy (anonymous):

Simplify. square root 108 minus 3 times the square root of 8 minus the square root 27

OpenStudy (kc_kennylau):

Do you mean \(\sqrt{180}-3\sqrt{8}-\sqrt{27}\)?

OpenStudy (anonymous):

10.39-12-5.19= -6.8

OpenStudy (kc_kennylau):

Simplify. Not plug it into a calculator.